# SpiroLLM: Finetuning Pretrained LLMs to Understand Spirogram Time Series with Clinical Validation in COPD Reporting

### v2

@@ Acknowledgements
SH, YZ, and JS conceptualized the study, acquired the funding, supervised, and administered the project. The methodology was developed by SM with contributions from YL, SH, and YZ. SM was responsible for the software development, formal analysis, and visualization. YL contributed to the formal analysis. SC and XH performed the investigation and were responsible for data curation. The results were validated by SC, XH, and SG. SM wrote the original draft of the manuscript. SH, YZ, and JS were major contributors in reviewing and editing the manuscript. All authors read and approved the final manuscript.
@@
The authors declare no competing interests.
@@ Appendix F Expert Evaluation Results
Figure 6: Evaluation by Pulmonary Function Experts.
Figure 7: Evaluation by Pulmonary Function Experts.
Figure 8: Evaluation by Pulmonary Function Experts.
Figure 9: Evaluation by Pulmonary Function Experts.

### v3

@@ Acknowledgements
SH, YZ, and JS conceptualized the study, acquired the funding, supervised, and administered the project. The methodology was developed by SM with contributions from YL, SH, and YZ. SM was responsible for the software development and formal analysis. SM and XX were responsible for visualization. YL contributed to the formal analysis. SC and XH performed the investigation and were responsible for data curation. The results were validated by SC, XH, and SG. SM wrote the original draft of the manuscript. SH, YZ, and JS were major contributors in reviewing and editing the manuscript. All authors read and approved the final manuscript.
@@
The authors have declared that no competing interests exist.